OpenStudy (anonymous):

xy+wz= w=3 x=6 y=5 and z=9

OpenStudy (anonymous):

Substitute the numbers given for the appropriate letters, then use order of operations to simplify.

OpenStudy (texaschic101):

xy + wz .....w = 3, x = 6, y = 5, z = 9 now we sub (6)(5) + (3)(9) = 30 + 27 = 57

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

thats all

OpenStudy (arabpride):

Please close this post if you're done. Thanks! :D